# GenAI vs. Human Creators: Procurement Mechanism Design in Two-/Three-Layer Markets

### v1

@@ Abstract
With the rapid advancement of generative AI (GenAI), mechanism design adapted to its unique characteristics poses new theoretical and practical challenges. Unlike traditional goods, content from one domain can enhance the training and performance of GenAI models in other domains. For example, OpenAI’s video generation model Sora (liu2024sora) relies heavily on image data to improve video generation quality. In this work, we study nonlinear procurement mechanism design under data transferability, where online platforms employ both human creators and GenAI to satisfy cross-domain content demand. We propose optimal mechanisms that maximize either platform revenue or social welfare and identify the specific properties of GenAI that make such high-dimensional design problems tractable. Our analysis further reveals which domains face stronger competitive pressure and which tend to experience overproduction. Moreover, the growing role of data intermediaries, including labeling companies such as Scale AI and creator organizations such as The Wall Street Journal, introduces a third layer into the traditional platform–creator structure. We show that this three-layer market can result in a lose-lose outcome, reducing both platform revenue and social welfare, as large pre-signed contracts distort creators’ incentives and lead to inefficiencies in the data market. These findings suggest a need for government regulation of the GenAI data ecosystem, and our theoretical insights are further supported by numerical simulations.
@@ 1 Introduction
Large models have entered the era of multimodality (yin2023survey). Modern commercial systems such as GPT-4 (achiam2023gpt), Gemini (team2023gemini), and Claude (Claude3ModelCard) are trained on diverse types of data, including text, images, audio, and video. These modalities are not isolated; transferability among them is well established both theoretically and empirically through transfer learning (howard2018universal). Just as film directors may draw inspiration from Shakespeare’s works, image datasets can enhance the performance of video generation models. This interdependence links the valuation of data across domains into a unified economic problem and naturally raises the question:
@@
The transferability of data also raises important ethical concerns. Recent debates have centered on whether GenAI will cause certain professions to vanish, alongside growing discussions on the implications of artificial general intelligence (george2023chatgpt). We want to ask:
@@
The growing demand for data to train large models has driven many AI companies to purchase high-quality datasets. For example, OpenAI reportedly spent about \$250 million acquiring data from The Wall Street Journal (wsj2024openai). Industry forecasts further predict that the global generative AI market for content creation will expand from USD 15.2 billion in 2024 to USD 175.3 billion by 2033 (MarketUs2024). This surge in data demand can be traced back to the creation of ImageNet (deng2009imagenet), which catalyzed the modern data economy. The emergence of specialized data annotation companies such as Scale AI, Lionbridge, Aurora AI, and Amazon Mechanical Turk has since transformed the landscape, pushing the market from a two-layer structure toward a three-layer structure that includes professional data brokers. This evolution raises a key question:
@@ 1.1 Our Contributions
Recent studies have begun examining market equilibria that arise from competition between AI content generators and human creators (yao2024human). Our work is the first to approach this problem from the buyer’s perspective. We study online sharing platforms that rely on both human creators and GenAI for content production, situating the analysis within the broader framework of procurement mechanism design. Going beyond the classical single-dimensional setting (myerson1981optimal), we derive optimal mechanisms for five of six multi-dimensional environments and establish tight upper and lower bounds for revenue and social welfare in the remaining case. Our results show that although GenAI reduces the overall demand for human-created content, human creators remain indispensable from an economic standpoint. Their outputs not only satisfy subscriber preferences but also provide essential data for GenAI training. Furthermore, in domains with high data transferability, we find that overproduction is more likely to occur, reflecting the complex interactions between human and algorithmic production incentives.
@@
Finding an efficient market structure has been a long-standing goal in digital economics (liang2018survey; agarwal2019marketplace). For online platforms, a central question is whether to buy content directly from human creators as crowdsourcing, or through data companies. Our analysis shows that the latter, a three-layer market where the platform purchases data from a broker who sources it from humans, leads to a lose-lose outcome: both platform revenue and social welfare decline. The inefficiency stems from asymmetric information. Large, transparent contracts reveal the platform’s and broker’s valuations to creators, making price discrimination impossible and weakening incentives for efficient production. As a result, both welfare and revenue fall, underscoring the need for policy intervention.
@@ 1.2 Related Works
Our work is closely related to the literature on procurement mechanism design (myerson1981optimal; mcafee1987auctions; laffont1993theory), data pricing (jia2019towards; ghorbani2019data; schoch2022cs; ai2024instrumental), human vs. GenAI competition (esmaeili2024strategize; yao2024human; fish2024algorithmic; duetting2024mechanism) and multi-layer markets (fallah2024three; xu2020data). We generalize physical goods procurement to transferable digital content and analyze the ensuing human–GenAI competition across multi-layer markets. A thorough comparison and additional related work are provided in Section˜A.1.
@@ 2.1 A Model of Content Procurement in the GenAI Era
An online platform features $K$ domains (e.g., sports, pop music, knowledge sharing), denoted by $K]={1,2,\ldots,K}$. To serve its user base, the platform requires $d_{k}\in\mathbb{R}$ units of content in each domain $k\in[K]$, produced by human creators and/or GenAIs. We assume that each human creator specializes in a single domain [^1] Although some creators may produce content across domains, most major creators focus on one area in practice; see Figure 1 for examples.. For instance, Dream has attracted over 30 million YouTube subscribers through Minecraft challenge videos. For analytical convenience, each domain $k$ is represented by a single creator who incurs a cost $c_{k}x_{k}^{\rho_{k}}$ to produce $x_{k}$ units of content, following standard models of the creator economy (hu2023incentivizing; jagadeesan2024supply; yao2024human). We interpret $x_{k}$ as a calibrated quantity, meaning that production is weighted by quality metrics such as click-through rate or content accuracy (radlinski2008does). The parameter $c_{k}$ captures the creator’s production efficiency, which is private information, while $\rho_{k}$ reflects the population-level growth speed of the production cost and is assumed to be publicly known, as it can be inferred from market data. This representative-creator assumption is without loss of generality: when multiple creators exist in the same domain with heterogeneous cost parameters $c_{k}$, they can be aggregated into a single “meta” creator with an effective cost coefficient $c_{k}$ (see Section˜C.3 for details). Consistent with prior work, we assume $\rho_{k}\geq 1$ to reflect increasing marginal costs (marshall2009principles). This captures the empirical observation that creative ideas become progressively harder to generate over given time, resulting in a convex production cost function.
@@
To meet its demand requirements $(d_{1},\ldots,d_{K})$, the platform (e.g., a video-sharing site such as YouTube) asks human creators to report their private cost parameters $c=(c_{1},\ldots,c_{K})$. To account for possible misreporting, let $\widehat{c}=(\widehat{c}_{1},\ldots,\widehat{c}_{K})$ denote the reported costs. The mechanism then proceeds according to the standard timeline in mechanism design (hart1988contract; salant1989inducing):
@@
While procurement auctions have been extensively studied in prior work (laffont1993theory), our setting differs by incorporating the platform’s ability to use GenAI to augment content creation at negligible cost. Specifically, we assume that the platform can employ GenAI, together with human-created content, to partially meet demand with minimal additional expense relative to human creation. Motivated by the scaling laws of GenAI’s creative capacity (kaplan2020scaling; yao2024human), we model this process as follows: $x_{i}$ units of content in domain $i$ enable GenAI to generate $\mu_{ik}x_{i}^{\gamma_{ik}}$ units of essentially new content in domain $k$, where $\mu_{ik}$ captures the transferability from domain $i$ to domain $k$, and $\gamma_{ik}$ represents the capability of employed GenAI tool. This formulation highlights three key distinctions between GenAI, human creators, and traditional machines. First, GenAI exhibits transferability, allowing knowledge from one domain to benefit others with varying efficiency. Second, its marginal cost of production is negligible compared with human effort. Third, GenAI’s productivity depends on human creation, unlike traditional machines. The power-law relationship between model capability and training data is well documented in the GenAI literature (kaplan2020scaling), and we assume that these capability parameters are known, as they can be empirically estimated (alabdulmohsin2022revisiting; goyal2024scaling; lin2024selecting). Finally, as is standard, we restrict $\gamma_{ik}\in[0,1]$ to reflect the diminishing efficiency of data. Accordingly, given reported costs $\widehat{c}$, the platform’s design must satisfy the following demand constraints:
@@
The revelation principle (roughgarden2010algorithmic) implies that, without loss of generality, we can restrict attention to incentive-compatible (IC) mechanisms in which truthful reporting constitutes a Nash equilibrium (myerson1986multistage). Accordingly, the incentive-compatibility condition is given by
@@
where $c_{-k}$ is $c$ except the $k$-th entry. The expectation is taken over domains $i\neq k$. The nonlinearity of the cost function hinders the application of existing methods (myerson1981optimal). We first focus on the Nash equilibrium (nash1950equilibrium), assuming no collusion, so it suffices to consider one-shot deviations. We will later extend the analysis to more complex settings involving creator union and data brokers. An interesting future direction is to extend it to correlated equilibrium and coarse correlated equilibrium (aumann1987correlated) with limited signals.
@@
In practice, platforms typically pursue one of two objectives. Some aim to maximize revenue, or equivalently, minimize total cost. Others, such as large technology firms like Google (google_social_impact) and Meta (meta_social_impact), emphasize social welfare, defined as the aggregate utility of all market participants, or equivalently, the minimization of creators’ total costs, since monetary transfers do not affect overall welfare. These considerations motivate the two optimization objectives:
@@
To avoid lengthy discussions of corner cases, we impose standard regularity assumptions on the above distributions, as commonly adopted in both the economics and machine learning literature (myerson1981optimal; wang2019optimism; ai2022reinforcement).
@@
Assumption˜2.2 can always be ensured through distribution truncation, while Assumption˜2.3 is a standard and widely popular assumption in economics (kleiber2003statistical; bagnoli2006log; wang2024relyingmetricsevaluatedagents). These assumptions are mild and hold for many commonly used distributions, such as the truncated Gaussian and the uniform distribution (golrezaei2019dynamic).
@@ 2.2 The Optimal Mechanism $\mathcal{M}_{1}$ for Revenue-Maximizing Procurement
While Theorem˜2.1 is formulated in a single-parameter mechanism design setting, the main challenge in its proof arises from the nonlinearity of the designer’s allocation constraint, namely $y_{k}^{1/\rho_{k}}+\sum_{i=1}^{K}\mu_{ik}y_{i}^{\gamma_{ik}/\rho_{i}}\geq d_{k}$ in Algorithm˜1. This constraint, induced by the nonlinear cost function and data transferability, prevents the direct application of Myerson’s standard analysis. We address this difficulty in two steps. First, by applying a nonlinear transformation, we replace the search for the optimal allocation $x(\cdot)$ with that for $y(\cdot)$, a monotonic transformation of $x(\cdot)$, which linearizes the cost term. Second, we show that the nonlinear constraints introduced by GenAI still preserve convexity. Together, these steps reformulate Problem 1 into a convex optimization problem, for which we establish the existence of an efficient solution algorithm.
@@
Randomized payments can further reduce the computational complexity to $\mathcal{O}(1/\epsilon^{2})$ and $\mathcal{O}(1/\epsilon)$, as detailed in Proposition˜C.1 of Section˜C.1. These results demonstrate that online platforms can efficiently maximize revenue even when incorporating GenAI into content creation. This extends the scope of procurement mechanism design and provides a theoretical foundation for pricing in the rapidly evolving GenAI economy.
@@ 2.4 Creator Union and Optimal Procurement Mechanism Design
Recently, OpenAI and The Wall Street Journal reached a data agreement valued at up to \$250 million (wsj2024openai). At the same time, data-selling companies such as Scale AI have experienced rapidly rising valuations, driven by the growing demand for high-quality data to fine-tune large language models (scale2023openai). Economically, these intermediaries can be viewed as aggregating data from human creators and repackaging it for sale, similar to a creator union. We defer the analysis of the case in which such intermediaries secure contracts before sourcing data from creators to Section˜3. This raises a key question: when a creator union appears, represented by a single agent producing content across all $K$ domains, how will the revenue of data-obsessing platforms and social welfare be affected?
We retain the setting from Section˜2.1, except that the representative creator now has a cost vector $c=(c_{1},\ldots,c_{K})$. We also let $a=(a_{1},\ldots,a_{K})$ and $b=(b_{1},\ldots,b_{K})$. Unlike before, one-shot deviations are insufficient to capture strategic behavior, as the creator may misreport multiple entries simultaneously. This possibility of collusion complicates the mechanism design problem by introducing correlations, analogous to the distinction between Nash and correlated equilibria (gilboa1989nash).

### v2

@@ Abstract
With the rapid advancement of generative AI (GenAI), mechanism design adapted to its unique characteristics poses new theoretical and practical challenges. Unlike traditional goods, content from one domain can enhance the training and performance of GenAI models in other domains. For example, OpenAI’s video generation model Sora (Liu et al., 2024b) relies heavily on image data to improve video generation quality. In this work, we study nonlinear procurement mechanism design under data transferability, where online platforms employ both human creators and GenAI to satisfy cross-domain content demand. We propose optimal mechanisms that maximize either platform revenue or social welfare and identify the specific properties of GenAI that make such high-dimensional design problems tractable. Our analysis further reveals which domains face stronger competitive pressure and which tend to experience overproduction. Moreover, the growing role of data intermediaries, including labeling companies such as Scale AI and creator organizations such as The Wall Street Journal, introduces a third layer into the traditional platform–creator structure. We show that this three-layer market can result in a lose-lose outcome, reducing both platform revenue and social welfare, as large pre-signed contracts distort creators’ incentives and lead to inefficiencies in the data market. These findings suggest a need for government regulation of the GenAI data ecosystem, and our theoretical insights are further supported by numerical simulations.
@@ 1 Introduction
Large models have entered the era of multimodality (Yin et al., 2023). Modern commercial systems such as GPT-4 (Achiam et al., 2023), Gemini (Team et al., 2023), and Claude (Anthropic, 2024) are trained on diverse types of data, including text, images, audio, and video. These modalities are not isolated; transferability among them is well established both theoretically and empirically through transfer learning (Howard and Ruder, 2018). Just as film directors may draw inspiration from Shakespeare’s works, image datasets can enhance the performance of video generation models. This interdependence links the valuation of data across domains into a unified economic problem and naturally raises the question:
@@
The transferability of data also raises important ethical concerns. Recent debates have centered on whether GenAI will cause certain professions to vanish, alongside growing discussions on the implications of artificial general intelligence (George et al., 2023). We want to ask:
@@
The growing demand for data to train large models has driven many AI companies to purchase high-quality datasets. For example, OpenAI reportedly spent about \$250 million acquiring data from The Wall Street Journal (WSJ, 2024). Industry forecasts further predict that the global generative AI market for content creation will expand from USD 15.2 billion in 2024 to USD 175.3 billion by 2033 (Market.us, 2024). This surge in data demand can be traced back to the creation of ImageNet (Deng et al., 2009), which catalyzed the modern data economy. The emergence of specialized data annotation companies such as Scale AI, Lionbridge, Aurora AI, and Amazon Mechanical Turk has since transformed the landscape, pushing the market from a two-layer structure toward a three-layer structure that includes professional data brokers. This evolution raises a key question:
@@ 1.1 Our Contributions
Recent studies have begun examining market equilibria that arise from competition between AI content generators and human creators (Yao et al., 2024). Our work is the first to approach this problem from the buyer’s perspective. We study online sharing platforms that rely on both human creators and GenAI for content production, situating the analysis within the broader framework of procurement mechanism design. Going beyond the classical single-dimensional setting (Myerson, 1981), we derive optimal mechanisms for five of six multi-dimensional environments and establish tight upper and lower bounds for revenue and social welfare in the remaining case. Our results show that although GenAI reduces the overall demand for human-created content, human creators remain indispensable from an economic standpoint. Their outputs not only satisfy subscriber preferences but also provide essential data for GenAI training. Furthermore, in domains with high data transferability, we find that overproduction is more likely to occur, reflecting the complex interactions between human and algorithmic production incentives.
@@
Finding an efficient market structure has been a long-standing goal in digital economics (Liang et al., 2018; Agarwal et al., 2019). For online platforms, a central question is whether to buy content directly from human creators as crowdsourcing, or through data companies. Our analysis shows that the latter, a three-layer market where the platform purchases data from a broker who sources it from humans, leads to a lose-lose outcome: both platform revenue and social welfare decline. The inefficiency stems from asymmetric information. Large, transparent contracts reveal the platform’s and broker’s valuations to creators, making price discrimination impossible and weakening incentives for efficient production. As a result, both welfare and revenue fall, underscoring the need for policy intervention.
@@ 1.2 Related Works
Our work is closely related to the literature on procurement mechanism design (Myerson, 1981; McAfee and McMillan, 1987; Laffont and Tirole, 1993), data pricing (Jia et al., 2019; Ghorbani and Zou, 2019; Schoch et al., 2022; Ai et al., 2024), human vs. GenAI competition (Esmaeili et al., 2024; Yao et al., 2024; Fish et al., 2024; Duetting et al., 2024) and multi-layer markets (Fallah et al., 2024; Xu et al., 2020). We generalize physical goods procurement to transferable digital content and analyze the ensuing human–GenAI competition across multi-layer markets. A thorough comparison and additional related work are provided in Section˜A.1.
@@ 2.1 A Model of Content Procurement in the GenAI Era
An online platform features $K$ domains (e.g., sports, pop music, knowledge sharing), denoted by $K]={1,2,\ldots,K}$. To serve its user base, the platform requires $d_{k}\in\mathbb{R}$ units of content in each domain $k\in[K]$, produced by human creators and/or GenAIs. We assume that each human creator specializes in a single domain [^1] Although some creators may produce content across domains, most major creators focus on one area in practice; see Figure 1 for examples.. For instance, @dream has attracted over 30 million YouTube subscribers through Minecraft challenge videos. For analytical convenience, each domain $k$ is represented by a single creator who incurs a cost $c_{k}x_{k}^{\rho_{k}}$ to produce $x_{k}$ units of content, following standard models of the creator economy (Hu et al., 2023; Jagadeesan et al., 2024; Yao et al., 2024). We interpret $x_{k}$ as a calibrated quantity, meaning that production is weighted by quality metrics such as click-through rate or content accuracy (Radlinski et al., 2008). The parameter $c_{k}$ captures the creator’s production efficiency, which is private information, while $\rho_{k}$ reflects the population-level growth speed of the production cost and is assumed to be publicly known, as it can be inferred from market data. This representative-creator assumption is without loss of generality: when multiple creators exist in the same domain with heterogeneous cost parameters $c_{k}$, they can be aggregated into a single “meta” creator with an effective cost coefficient $c_{k}$ (see Section˜C.4 for details). Consistent with prior work, we assume $\rho_{k}\geq 1$ to reflect increasing marginal costs (Marshall, 2009). This captures the empirical observation that creative ideas become progressively harder to generate over given time, resulting in a convex production cost function.
@@
To meet its demand requirements $(d_{1},\ldots,d_{K})$, the platform (e.g., a video-sharing site such as YouTube) asks human creators to report their private cost parameters $c=(c_{1},\ldots,c_{K})$. To account for possible misreporting, let $\widehat{c}=(\widehat{c}_{1},\ldots,\widehat{c}_{K})$ denote the reported costs. The mechanism then proceeds according to the standard timeline in mechanism design (Hart and Tirole, 1988; Salant, 1989):
@@
While procurement auctions have been extensively studied in prior work (Laffont and Tirole, 1993), our setting differs by incorporating the platform’s ability to use GenAI to augment content creation at negligible cost (We discuss the non-negligible case in Section˜C.1). Specifically, we assume that the platform can employ GenAI, together with human-created content, to partially meet demand with minimal additional expense relative to human creation. Motivated by the scaling laws of GenAI’s creative capacity (Kaplan et al., 2020; Yao et al., 2024), we model this process as follows: $x_{i}$ units of content in domain $i$ enable GenAI to generate $\mu_{ik}x_{i}^{\gamma_{ik}}$ units of essentially new content in domain $k$, where $\mu_{ik}$ captures the transferability from domain $i$ to domain $k$, and $\gamma_{ik}$ represents the capability of employed GenAI tool. Yao et al. (2024) adopts a similar assumption to analyze the symbiosis or conflict between GenAI and human creators. This formulation highlights three key distinctions between GenAI, human creators, and traditional machines. First, GenAI exhibits transferability, allowing knowledge from one domain to benefit others with varying efficiency. Second, its marginal cost of production is negligible compared with human effort. Third, GenAI’s productivity depends on human creation, unlike traditional machines. The power-law relationship between model capability and training data is well documented in the GenAI literature (Kaplan et al., 2020), and we assume that these capability parameters are known, as they can be empirically estimated (Alabdulmohsin et al., 2022; Goyal et al., 2024; Lin et al., 2024). Finally, as is standard, we restrict $\gamma_{ik}\in[0,1]$ to reflect the diminishing efficiency of data. Accordingly, given reported costs $\widehat{c}$, the platform’s design must satisfy the following demand constraints:
@@
The revelation principle (Roughgarden, 2010) implies that, without loss of generality, we can restrict attention to incentive-compatible (IC) mechanisms in which truthful reporting constitutes a Nash equilibrium (Myerson, 1986). In Section˜A.2, we provide illustrations about the economic background for these and other relevant terms. Accordingly, the incentive-compatibility condition is given by
@@
where $c_{-k}$ is $c$ except the $k$-th entry. The expectation is taken over domains $i\neq k$. The nonlinearity of the cost function hinders the application of existing methods (Myerson, 1981). We first focus on the Nash equilibrium (Nash Jr, 1950), assuming no collusion, so it suffices to consider one-shot deviations. We will later extend the analysis to more complex settings involving creator union and data brokers. An interesting future direction is to extend it to correlated equilibrium and coarse correlated equilibrium (Aumann, 1987) with limited signals.
@@
In practice, platforms typically pursue one of two objectives. Some aim to maximize revenue, or equivalently, minimize total cost. Others, such as large technology firms like Google (Google, 2024) and Meta (Meta, 2024), emphasize social welfare, defined as the aggregate utility of all market participants, or equivalently, the minimization of creators’ total costs, since monetary transfers do not affect overall welfare. These considerations motivate the two optimization objectives:
@@
To avoid lengthy discussions of corner cases, we impose standard regularity assumptions on the above distributions, as commonly adopted in both the economics and machine learning literature (Myerson, 1981; Wang et al., 2019; Ai et al., 2022).
@@
Assumption˜2.2 can always be ensured through distribution truncation, while Assumption˜2.3 is a standard and widely popular assumption in economics (Kleiber, 2003; Bagnoli and Bergstrom, 2006; Wang et al., 2024). These assumptions are mild and hold for many commonly used distributions, such as the truncated Gaussian and the uniform distribution (Golrezaei et al., 2019).
@@ 2.2 The Optimal Mechanism $\mathcal{M}_{1}$ for Revenue-Maximizing Procurement
While Theorem˜2.1 is formulated in a single-parameter mechanism design setting, the main challenge in its proof arises from the nonlinearity of the designer’s allocation constraint, namely $y_{k}^{1/\rho_{k}}+\sum_{i=1}^{K}\mu_{ik}y_{i}^{\gamma_{ik}/\rho_{i}}\geq d_{k}$ in Algorithm˜1. This constraint, induced by the nonlinear cost function and data transferability, prevents the direct application of the standard analysis by Myerson (1981). We address this difficulty in two steps. First, by applying a nonlinear transformation, we replace the search for the optimal allocation $x(\cdot)$ with that for $y(\cdot)$, a monotonic transformation of $x(\cdot)$, which linearizes the cost term. Second, we show that the nonlinear constraints introduced by GenAI still preserve convexity. Together, these steps reformulate Problem 1 into a convex optimization problem, for which we establish the existence of an efficient solution algorithm.
@@
Randomized payments can further reduce the computational complexity to $\mathcal{O}(1/\epsilon^{2})$ and $\mathcal{O}(1/\epsilon)$, as detailed in Proposition˜C.1 of Section˜C.2. These results demonstrate that online platforms can efficiently maximize revenue even when incorporating GenAI into content creation. This extends the scope of procurement mechanism design and provides a theoretical foundation for pricing in the rapidly evolving GenAI economy.
@@ 2.4 Creator Union and Optimal Procurement Mechanism Design
Recently, OpenAI and The Wall Street Journal reached a data agreement valued at up to \$250 million (WSJ, 2024). At the same time, data-selling companies such as Scale AI have experienced rapidly rising valuations, driven by the growing demand for high-quality data to fine-tune large language models (The Scale Team, 2023). Economically, these intermediaries can be viewed as aggregating data from human creators and repackaging it for sale, similar to a creator union. We defer the analysis of the case in which such intermediaries secure contracts before sourcing data from creators to Section˜3. This raises a key question: when a creator union appears, represented by a single agent producing content across all $K$ domains, how will the revenue of data-obsessing platforms and social welfare be affected?
We retain the setting from Section˜2.1, except that the representative creator now has a cost vector $c=(c_{1},\ldots,c_{K})$. We also let $a=(a_{1},\ldots,a_{K})$ and $b=(b_{1},\ldots,b_{K})$. Unlike before, one-shot deviations are insufficient to capture strategic behavior, as the creator may misreport multiple entries simultaneously. This possibility of collusion complicates the mechanism design problem by introducing correlations, analogous to the distinction between Nash and correlated equilibria (Gilboa and Zemel, 1989).